OpenStudy (anonymous):

the ratio between the circumference of two circles is 4:9,find the ratio between their areas.

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):



OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{ 2\pi ~oa=4k\\~\\ 2\pi ~ob=9k\\~\\ \dfrac{oa}{ob}=\dfrac{4}{9}\\~\\ \dfrac{oa^2}{ob^2}=\dfrac{16}{81}\\~\\ \dfrac{\pi ~oa^2}{\pi ~ob^2}=\dfrac{16}{81}\\~\\ }\end{align}\)

OpenStudy (anonymous):

?

OpenStudy (mathmath333):

what is formula for area of circle ?

OpenStudy (anonymous):

pi^r^2

OpenStudy (mathmath333):

it is \(\pi~r^2\) right

OpenStudy (anonymous):

yes

OpenStudy (mathmath333):

what is formula for circumference of circle

OpenStudy (anonymous):

2^pi^r

OpenStudy (mathmath333):

u mean \(2\pi r\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes

OpenStudy (mathmath333):

so \(\dfrac{2\pi r_1}{2\pi r_2}=\dfrac{4}{9}\) is given , u agree

OpenStudy (mathmath333):

so \(\dfrac{ r_1}{r_2}=\dfrac{4}{9}\) u agree

OpenStudy (anonymous):

yep

OpenStudy (mathmath333):

so \(\dfrac{ r_1^2}{r_2^2}=\dfrac{16}{81}\) u stilll agree

OpenStudy (anonymous):

yes

OpenStudy (mathmath333):

ok \(\dfrac{ \pi r_1^2}{\pi r_2^2}=\dfrac{16}{81}\) still yes

OpenStudy (anonymous):

yes

OpenStudy (mathmath333):

ok \(\dfrac{A_1}{A_2}=\dfrac{\pi r_1^2}{\pi r_2^2}\) still yes

OpenStudy (anonymous):

yes

OpenStudy (mathmath333):

what is \(A_1\) and \(A_2\)

OpenStudy (anonymous):

16/81?

OpenStudy (mathmath333):

yes

OpenStudy (anonymous):

now?

OpenStudy (mathmath333):

these are the ratio of the areas,u wanted to find .lol

OpenStudy (anonymous):

one more question

OpenStudy (mathmath333):

ok

OpenStudy (anonymous):

the wheel of a cart is making 5 revolutions per sec . if the diameter of the wheel is 84cm find the speed in kn/hr give ur answer in the correct nearest km

OpenStudy (mathmath333):

u mean speed in km/hr ?

OpenStudy (anonymous):

yes that was a typo i need it in km/hr.

OpenStudy (anonymous):

help?

OpenStudy (mathmath333):

what is circumference of the wheel

OpenStudy (anonymous):

44 x 6?

OpenStudy (mathmath333):

\(\huge C=2\pi r=\pi d\)

OpenStudy (anonymous):

orry its 22 x 12

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

264cm?

OpenStudy (mathmath333):

\(C=\pi d=\dfrac{22}{7}\times 84 \)

OpenStudy (mathmath333):

yes

OpenStudy (anonymous):

now?

OpenStudy (mathmath333):

so one revolution means it travels 264 cm , u agree

OpenStudy (anonymous):

yes

OpenStudy (mathmath333):

so how much distance it travels in one second

OpenStudy (anonymous):

264cm

OpenStudy (mathmath333):

the wheel of a cart is making 5 revolutions per sec ^ given in question

OpenStudy (mathmath333):

1 revolution =264 cm 5 revolution =?

OpenStudy (anonymous):

1320?

OpenStudy (anonymous):

1320cm?

OpenStudy (mathmath333):

yes but it is given in unit \(\dfrac{cm}{sec}\) u want it in \(\dfrac{km}{hr}\)

OpenStudy (anonymous):

0.0132km

OpenStudy (mathmath333):

1cm =...km ?

OpenStudy (mathmath333):

yes

OpenStudy (anonymous):

thats the answer?

OpenStudy (mathmath333):

1 sec= .. . hr?

OpenStudy (anonymous):

3600

OpenStudy (mathmath333):

u mean 1 sec =3600 hrs ??

OpenStudy (anonymous):

i meant 1 hr = 3600 secs

OpenStudy (mathmath333):

so 1 sec = ..hrs?

OpenStudy (anonymous):

idk

OpenStudy (mathmath333):

\(1~sec=\dfrac{1}{3600} hrs\) u agree

OpenStudy (mathmath333):

\(1~hr =3600~seconds\) divide both sides by \(3600\) what will u get

OpenStudy (anonymous):

0.00027?

OpenStudy (anonymous):

0.00027hr?

OpenStudy (mathmath333):

yes

OpenStudy (mathmath333):

so \(\dfrac{0.0132}{0.00027}\)=

OpenStudy (anonymous):

idk

OpenStudy (mathmath333):

so it is \(\approx\dfrac{48~km}{hr}\)

OpenStudy (anonymous):

thx,i hv more q's

Directrix (directrix):

Another way to look at the first question: Theorem: If two plane figures are similar, the square of the scale factor of the two figures is equal to the ratio of any two corresponding area measurements of the figures. All circles are similar. Circumference is linear. (4/9)^2 = ratio of areas 16/81 = ratio of areas of the two circles.